OpenStudy (anonymous):

problem: give the exact value of sec(pi/3). so i understand that sec=1/cos, but every time i try to solve it i get it wrong. can someone explain?

OpenStudy (***[isuru]***):

hey... ur point seems confusing but... cos (pi/3) = 1/2 and hence sec (pi/3) = 1/cos(pi/3) = 1/ (1/2) = 2 i don't know whether this is the help u seek out.. but if u could explain ur situation a bit more it would be helpful.. :)

OpenStudy (anonymous):

for some reason, every time i type cos(pi/3) into my calculator is comes out as .9976035, but i think i was just typing it in wrong. thanks for the help!

OpenStudy (***[isuru]***):

u r welcome :)